### v3

# UFGraphFR: Graph Federation Recommendation System based on User Text description features

Abstract
Federated learning has become a key paradigm that utilizes the "data available but invisible" feature to protect user privacy, especially in the artificial intelligence field where user privacy is highly valued. Federated learning enables users to collaboratively train models without sharing raw data. Motivated by this, federated recommendation systems offer a promising architecture that balances user privacy with recommendation accuracy through distributed collaborative learning. However, existing federated recommendation methods often neglect the underlying semantic or behavioral relationships between users during parameter aggregation, which limits their recommendation effectiveness. To overcome this limitation, graph-based federated recommendation systems have been proposed to leverage neighborhood information. Yet, conventional graph construction methods usually require access to raw user data or explicit social links, which contradicts the strict privacy requirements of federated learning. In this paper, we propose UFGraphFR, a novel model designed to effectively capture user relationships for constructing user graphs while preserving the confidentiality of users’ raw data. (1) To better utilize users’ textual feature descriptions under the hypothesis that users with similar textual profiles exhibit analogous preferences, we propose a Joint-Embedding Layer architecture designed to encode structured user features into fixed-dimensional vector representations. (2) To prevent raw data exposure during temporal pattern extraction, we introduce a Transformer-based sequence encoder with differential privacy guarantees, which employs multi-head self-attention mechanisms to model long-range dependencies in user behavioral sequences through on-device processing. These embeddings are then used to construct a semantic user graph for server-side aggregation via a lightweight graph neural architecture. UFGraphFR ensures privacy-preserving, personalized recommendations while maintaining model robustness and convergence efficiency under heterogeneous conditions. The code is available at: https://github.com/trueWangSyutung/UFGraphFR.
keywords:
Federated Learning;Federated Recommendation System;Recommendation
journal: Arxiv
[…]
[label1]organization=School of Computer Science and Engineering, addressline=Tianjin University of Technology, city=Tianjin, postcode=300384, state=Tianjin, country=China
{highlights}
Proposed a joint-embedding layer that converts structured user data into privacy-preserving text-based representations.
Developed a differentially private semantic graph construction method to model user correlations in federated learning.
Demonstrated superior performance over both federated and centralized baselines across multiple real-world datasets.
Validated framework effectiveness with different pre-trained language model embeddings.
1 Introduction
In the era of big data and privacy computing, intelligent recommendation systems have become indispensable infrastructure for modern digital services, enabling personalized movie suggestions, targeted food delivery recommendations, and even emotion-aware content curation through user comment analysis[1]. However, the conventional centralized paradigm that aggregates raw user data on servers poses critical privacy risks, as evidenced by growing user reluctance to share personal information with service providers[2]. In recent years, with the proposal of federated learning, which is a privacy-preserving framework that aims to better protect users’ privacy[3]. More and more research directions have started to apply this framework. Its emergence has catalyzed its application in the field of recommender systems, allowing the training of complete recommender system models without the raw data leaving the client device locally[4] .
Despite significant progress, there are two fundamental limitations of existing joint recommendation methods: (1). They model users as isolated entities mainly through local training goals, ignoring the inherent social correlations and behavioral dependencies among user groups[5], such as belonging to the same Beijing area, belonging to the same male gender, etc. This oversimplification contradicts real-world observations, where users with common interests often exhibit collaborative consumption patterns. (2). Despite the remarkable success of graph-enhanced recommender systems in centralized environments through neighborhood information aggregation, their direct application in federated learning is still prohibited due to privacy violations resulting from explicit user relationship disclosure[6] Combining the problems of the above two points, a graph augmentation based federated recommender system is proposed [7, 8]. (3). Current approaches mainly rely on Users’ ID-based feature representations[9], and such recommender systems perform poorly in the cold-start problem of models. Therefore, semantics can be utilized to enhance the cold start of the model. For example, two users can establish an association by the similarity between their textual features.
In this paper, we propose a novel framework for personalized federated recommendation, termed UFGraphFR (User Text description-based Feature Graph Federated Recommendation), which constructs a user relationship graph based on structured textual features without exposing private data. The proposed framework enables the server to aggregate user embeddings and generate personalized recommendations while complying with the privacy-preserving principle of federated learning. Specifically, our main contributions are summarized as follows:
  • 1.
    To better utilize user-structured attribute data, we introduce a joint-embedding layer that transforms user-side structured features into natural language descriptions, encodes them via a pre-trained language model into high-dimensional embeddings, and further projects them into compact representations via a lightweight linear transformation. The projection weights are shared with the server to guide semantic graph formation.
  • 2.
    To address the limitation that existing federated recommender systems ignore correlations between users, we propose a user graph construction method that builds a semantic user relationship graph on the server using differentially private embedding weights of user text descriptions. This allows the system to model collaborative preferences while maintaining strict data privacy.
  • 3.
    To validate the effectiveness of the proposed framework, we conduct extensive experiments on multiple real-world datasets. Results show that UFGraphFR achieves superior performance compared to existing federated and centralized baselines while preserving user privacy. And verify the effectiveness under different pre-trained language model embeddings.
2 Related Work
2.1 Federated Recommendation Systems
Federated Recommendation Systems (FedRS, as shown in Fig.1) is an important extension of Federated Learning (FL) in the field of recommendation, which realizes cross-client personalization modeling by protecting the user’s privacy. Personalized modeling, allows the user’s raw data to complete the training process of the recommender system without leaving the client while protecting the user’s privacy.
Some existing basic recommender system approaches (e.g., Matrix-Factorization-Based Recommender Systems[10], Neural-Collaborative-Filtering-Based Recommender Systems[11]) have been combined with the FL framework to form, for example, matrix factorization-based federated recommender systems FCF[12], FedMF[4], MetaMF[13], FedRecon[14], and neural collaborative filtering-based federated recommender system FedNCF[15].After this, more and more FedRSs were proposed. Zhang et al.[16] proposed a personalized FedRec framework, PFedRec, which removes user embedding and learns personalized scoring functions to capture user preferences. However, this framework ignores the correlation between users and does not respond well to the common preferences among similar users. To compensate for this shortcoming of correlation between users, Liu et al.[17] enhanced the federated recommender system by introducing the user’s social network. However, this framework exposes the user’s social network explicitly, which goes against the premise of protecting user privacy. Wu et al.[18] proposed FedPerGNN, an approach to capture user-item correlations using local map neural networks. At the same time, FedPerGNN also uses servers to find higher-order neighbors, thus providing more favorable information for local model training. However, this approach aligns historical interactions with the server leads to high computational overhead, and increases the risk of exposing users’ private information.
Therefore, correlation among users is a major pain point in federated recommender systems at this stage, and thus many researchers began to target FedRS frameworks with graph-guided mechanisms. Zhang et al.[1] proposed GPFedRec, a FedRS framework based on a graph-guided mechanism, GPFedRec designs a graph-guided aggregation mechanism to capture user preference correlations. Wang et al.[19] proposed CFedGR, a clustering-enhanced federated graph neural network recommender system, CFedGR proposes a clustering technique to enhance the user’s local first-order interaction subgraph. Wang et al.[20] proposed P4GCN, a Two-Party Graph Convolutional Network implementation of a vertical federated recommender system, which utilizes privacy-preserving Two-Party Graph Convolutional Networks (P4GCN) to improve the accuracy of recommendations without direct access to sensitive social information. So, the main direction of graph-guided FedRS at this stage is to construct user-user association graphs or user-item graphs without violating privacy and apply them to the whole federation training process.
However, FedRSs at this stage are focused on federated recommender systems based on the ID paradigm, which are limited in their semantic expressiveness, resulting in insufficient generalization ability and cold-start adaptation, as sufficient training data may not be available in such scenarios. Therefore, inspired by the success of Language Models (LMs) and their strong generalization ability, the potential of LMs can be exploited to enhance the capabilities of recommender systems. Luo et al[21]. proposed EasyRec, a centralized recommender system that employs a text-behavior alignment framework that combines contrastive learning with collaborative language model tuning to ensure tight alignment between the text-enhanced semantic space and collaborative behavioral information. Zhang et al[22]. proposed FedRS for IFedRec, which learns two sets of item representations by simultaneously utilizing item attributes and interaction records and designed an item representation alignment mechanism for aligning the two item representations within the FL framework and learning a network of meta-attributes on the server.
Inspired by the above, we combine graph-guided FedRS and LM-based FedRS to design a novel approach to build a user relationship graph, which utilizes users’ meta-attributes and pre-trained language models (PLMs) to construct a relationship graph between users. The aggregation of model parameters is achieved through this graph.
Another major issue in the Federated Learning process is how many clients are involved in the training -how the clients are selected, and the commonly used methods of client selection today are random selection and selection of all clients, of which random selection[23, 24, 25, 26, 27, 28, 29, 30] is the simplest method and is widely used in existing FedRSs, and is efficient in the case of a large number of clients being involved is very high. However, this method ignores the heterogeneity of clients, leading to suboptimal model convergence, as some clients may provide noisy or biased updates, which reduces the convergence of the model. Whereas Select All[31, 32, 33, 34, 35, 36] is an approach where all clients participate in each round of training, this approach has less efficient communication and is not suitable for devices with a large number of clients, but is often used in cross-platform FedRS approaches. So in our model, we use the full selection method.
Figure 1: a) shows the traditional centralized recommendation system. The user client first upload its own user characteristics and interactive data (with or without the consent of the user) to the server, and the server trains a recommendation model in the cloud to recommend suitable items for the user. b) shows a federated learning framework. Each client has a local model. The cloud sends the model parameters to the client in advance, and the client conducts local training without uploading private data to the cloud and uploals the trained model parameters to the server for aggregation.
2.2 Graph Learning-based Recommendation System
Recommendation systems for graph learning, as an emerging paradigm, are enhanced user (item) embeddings that learn by explicitly utilizing neighbor information in the graph structure. A common strategy is to integrate user-item interaction diagrams into a collaborative filtering framework. He et al. proposed a model called LightGCN[7], which applies graph convolutional networks to user-item interaction graphs to enrich representation learning for user preference prediction. We can also consider the adjacencies between all the items and form an item interaction graph. In sequence recommendation, it is also a good way to enhance learning through sequence diagrams[37]. With the rise of social networks, social recommendation based on social networks (the relationship network formed by using the social relationship between users) to enhance modeling gradually emerged[38]. In short, there is a lot of work going on to learn user embedding from user project interaction diagrams or social networks, or even to aggregate the two into a unified graph to enhance user representation. However, at a time when existing methods are increasingly focused on user privacy, this centralized recommendation system, which requires users to upload data to the cloud for training, is high-risk (as shown in Figure 1.a) because its centralized access to user data violates user privacy[1]. As a result, federal recommendations are gradually emerging, which combine privacy protection technologies to protect users’ private data from being leaked.
3 Preliminary
Federated Recommendation. The $U$ is user sets and the $I$ is item sets, respectively, let $r_{ui}$ be user-item interaction data between user $u$ and item $i$. Here is a recommendation system model $f$ for parameter $\theta$, which predicts $\hat{y_{ui}}=f(u,i|\theta)$ for users $u$ and $i$. On the central server, we represent the graph between all users with $\mathcal{G(U,E})$, where $\mathcal{U}$ represents the set of users and $\mathcal{E}$ represents the set of edges. At the same time, $\mathcal{A}={\{0,1\}}^{N\times N}$ represents its corresponding adjacency matrix form, and N represents the total number of users in the user set. $\mathcal{A}_{u}i=1$ indicates that user u and user i are associated.
For the above model, the purpose of the federated recommendation system is to predict user u’s preference for item i as $\hat{y_{ui}}=f(u,i|\theta^{*})$ , and the optimal model parameter at this time is $\theta^{*}$, as shown in formula (1).
$\theta^{*}=argmin_{\theta}=\sum_{i=1}^{N}\omega_{i}\mathcal{L}_{i}(\theta)$ (1)
Where $\mathcal{L}_{i}(\theta)$ is the loss of the local client participating in the training, and the parameter $\theta^{*}$ is learned by minimizing the local loss of all clients with the client weight $\omega_{i}$.
4 Methodology
In this section, we present the Graph Federation Recommendation System based on User Text description features(UFGraphFR). There are four steps in each communication round:
  • 1.
    Local Training: Clients initialize project embeddings using global parameters and train recommendation models with private interaction data.
  • 2.
    Parameter Uploading: Clients transmit user joint embedding weights and local item embeddings to the server.
  • 3.
    Graph Aggregation: The server constructs user relation graphs from text embeddings and aggregates parameters through graph convolution.
  • 4.
    Global Distribution: Updated global item embeddings are broadcast to all clients for next-round initialization.
4.1 Local Training
As shown in Figure 2, the client recommendation model consists of five major modules: 1) joint embedding layer (module parameter $\theta_{user}$) maps user structured attributes to vector space through natural language prompts templates, using a combination of a pre-trained language model initialization and a trainable linear layer; 2) item embedding layer (module parameter $\theta_{item}$) encodes user interaction item IDs as dense vectors; 3), textbfTemporal Transformer module (module parameter $\theta_{Transformer}$) models long-term interaction sequences through the mechanism of multi-head self-attention dependencies; 4) User Feature Refinement MLP (module parameter is $\theta_{\text{umlp}}$) extracts higher-order user representations through a three-layer fully-connected network; and 5) Predictive Scoring Function (module parameter is $\theta_{\text{score}}$) computes user-item embedding splicing based on final prediction results.
Figure 2: Client Model:The image shows our client model, with the yellow box on the left indicating the structure of the joint embedding layer (consisting of the prompt layer, pre-trained language model layer, and linear layer used to convert structured attributes into natural language). On the right is the Transformer Block used for sequence modeling
4.1.1 Joint Embedding Layer
We propose a joint embedding layer to deal with the user’s structured features, which contains a three-layer structure. These are the Prompts Output layer, the untrainable pre-trained language model (PLM) layer, and the trainable output layer.
Prompts Output Layer: Each user $u$, has a structured data $P_{u}$, as shown in Fig. 3. The role of Prompts Output Layer is to transform the structured data into textual language.
Figure 3: Prompts Output Layer: Structured data of the user on the left, textual description on the right
PLM Layer: After converting the structured features $P_{u}$ into natural language cues via $\Gamma(\dots)$, we obtain a string of user text features. This can in turn be transformed into d-dimensional high-dimensional vectors by freezing a pre-trained language model
$v_{u}=\text{LM}(\Gamma(\mathcal{P}_{u}))$ (2)
Where $\Gamma(\dots)$ denotes the work done by the Prompts Output Layer, i.e., transforming structured user attributes into natural language cue words, and $LM(\dots)$ denotes the transformation of the prompts into embedding vectors using a pre-trained language model.
However, the vector dimension of the embedded pre-trained natural language model depends on the chosen model (e.g., the embedding dimension of the USE model is 100, the embedding dimension of MiniLM-L6 is 384, and that of T5, TinyBERT, and LaBSE are all 768). To unify the fixed embedding dimensions, we add a trainable linear layer after the PLM layer to transform the high-dimensional embeddings into low-dimensional ones. Meanwhile, the weights of this linear layer are used to construct the user relationship graph. As shown in the formula (3), assuming that the vector dimension of the natural language model after embedding is $d_{1}$, then we need an embedding dimension of $d$.
$e_{u}=h(v_{u})=v_{u}W_{d_{1}\times d}+b$ (3)
Where $e_{u}$ is the vector embedded by the joint embedding layer, $W_{d_{1}\times d}$ is the weight matrix of this linear layer and $b$ is the bias of this linear layer.
4.1.2 Transformer Block
The Transformer Block in the local client model is a core component for capturing temporal dependencies in user-item interaction sequences. As illustrated in Figure 2, it consists of the following key operations:
Input Representation Given a user’s historical interaction sequence containing $L$ items, each item $i_{t}$ is embedded through the local item embedding layer $E_{item}$:
$e_{i_{t}}=E_{\text{item}}(i_{t})\in\mathbb{R}^{d_{\text{item}}}\quad(1\leq t\leq L)$ (4)
These embeddings are concatenated into a sequence matrix $X\in\mathbb{R}^{L\times d_{\text{item}}}$, which serves as the input to the Transformer Block.
Multi-Head Self-Attention Mechanism To capture the complex dependencies within user interaction sequences, we adopt a multi-head self-attention mechanism. Specifically, for each item embedding in the sequence, we compute contextualized representations by projecting the input matrix $X\in\mathbb{R}^{L\times d_{\text{item}}}$ into query ($Q$), key ($K$), and value ($V$) spaces through trainable linear transformations:
$Q=XW^{Q}\quad(W^{Q}\in\mathbb{R}^{d_{\text{item}}\times d_{h}})$ (5)
$K=XW^{K}\quad(W^{K}\in\mathbb{R}^{d_{\text{item}}\times d_{h}})$ (6)
$V=XW^{V}\quad(W^{V}\in\mathbb{R}^{d_{\text{item}}\times d_{h}})$ (7)
Then, for each attention head $k\in\{1,2,...,K\}$, we compute the scaled dot-product attention as:
$Z=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (8)
Here, $d_{h}=d_{\text{item}}/K$ ensures dimension alignment across multiple heads. This mechanism enables the model to attend to different parts of the sequence simultaneously, enriching the representation of long-range user preferences.
Feed-Forward Network To enhance the expressiveness of each position-wise representation, we apply a position-wise feed-forward network (FFN) after the self-attention layer. The FFN consists of two linear transformations with a ReLU activation in between:
$FFN(X)=\max(0,XW_{1}+b_{1})W_{2}+b_{2}$ (9)
This module refines the attended representations by introducing non-linearity and higher-level feature interactions, contributing to a more precise modeling of user-item interaction sequences.
4.1.3 Model Loss Functions
For the sake of generality, we discussed a typical scenario that relies only on implicit user-item interaction data for recommendations, i.e. if user $i$ interacts with item $j$, then $r_{ij}=1$; Otherwise, $r_{ij}=0$. No primitive features for secondary users (projects) are available. In this model, our loss function consists of two parts. The first part is the binary cross entropy Loss (BCE Loss) of prediction score and label, as shown in formula (10).
$\mathcal{L}_{1}(y,\hat{y})=-(ylog(\hat{y})+(1-y)log(1-\hat{y}))$ (10)
Where $y$ represents the user’s label, and $\hat{y}$ represents the predicted score output by the model. The second part is the regular term of the global item embedding and the user-specific item(sampling negative term) embedding, as shown in formula (11).
$\mathcal{R}(e_{global},e_{i}^{-})=Mean((e_{global}-e_{i}^{-})^{2})$ (11)
Where $e_{i}^{-}$ represents the embedding of sampling negative terms on the client. $e_{global}$ is the global embedded weight of the server-side aggregation. Based on the above formula, our total loss function is shown in formula (12).
$\mathcal{L}_{all}=\mathcal{L}_{1}(y,\hat{y})+\lambda\mathcal{R}(e_{global},e_{i}^{-})$ (12)
Where $\lambda$ is a hyperparameter, represents the regularization coefficient..
4.2 Federated Training
Figure 4: Service Model: First, it receives the parameters passed up from the client; second, it constructs the user relationship graph based on the parameters of the Jointed-Embedding layer; third, it learns the global item embedding; finally, it sends the global item embedding down to the client.
In the federated learning scenario, which is shown in Fig. 4, where we classify the parameters of the user model into two categories, and the model parameters are classified into two categories based on privacy sensitivity: globally shared parameters $\theta_{global}$ and local private parameters $\theta_{local}$. At each time, the user only uploads the globally shared parameters to the server for parameter aggregation. In our model, global parameters have joint embedding $\theta_{user}$ and item embedding $\theta_{item}$.
By alternately optimizing the local loss function and the global regular term Eq. (13), the model achieves cross-client knowledge migration while protecting the privacy of the original interaction data.
$\min_{\{\theta_{all,1}\dots\theta_{all,N}\}}\sum_{i=1}^{N}\mathcal{L}_{i}(\theta_{all,i})+\lambda\mathcal{R}_{i}(\theta_{global},\theta_{item,i})$ (13)
Where $\theta_{all,i}=\{\theta_{item,i},\theta_{user,i},\theta_{umlp,i},\theta_{score,i},\theta_{Transformer,i}\}$ is the recommended model parameter for the $i$ th client, and ri is the global user item embed weight aggregated on the server. $\mathcal{R}(\textperiodcentered,\textperiodcentered)$ is a regularization term used to constrain local item embedding similar to global user item embedding weights, where $\lambda$ is the regularization coefficient.
4.2.1 Build User Relaction Graph
Traditional federated learning frameworks (e.g., FedAvg[39]) use a parameter averaging aggregation strategy:
$\theta_{\text{global}}^{t+1}=\frac{1}{N}\sum_{i=1}^{N}\theta_{\text{item},i}^{t}$ (14)
This method treats all clients equally, but ignores group similarity in user preferences. Therefore, in our model, in order to further capture the correlation between users, we construct a dynamic relationship graph $G=(V,E)$ based on the joint embedding weights uploaded by clients on the server side, which follows the following process:
Step1: User Embedding Vectorization As described above when introducing the locally trained model, our proposed joint embedding layer consists of three parts, i.e., a cueing layer, an untrainable pre-trained language model embedding layer with embedding dimension $d_{1}$, and a trainable linear layer with output dimension $d_{2}$. Therefore, the parameter of the joint embedding layer is the embedding weight $W$ of the linear layer. It is a matrix of dimension $d_{1}\times d_{2}$, so the goal of the first step is to extend $W_{i}\in\mathbb{R}^{d_{1}\times d_{2}}$ to a vector of dimension $d_{1}\times d_{2}$ $w_{i}=\text{vec}(W_{i})\in\mathbb{R}^{d_{1}d_{2}}$.
Step2: Similarity computation constructed graph In the second step, we use cosine similarity as a similarity measure between two vectors $w_{i}$ and $w_{j}$ for users $i$ and $j$, and the similarity between customer $i$ and $j$ can be expressed as (15).
$\mathcal{S}_{ij}=\frac{w_{i}\textperiodcentered w_{j}}{||\vec{w}_{i}||||\vec{w}_{j}||}$ (15)
where $w_{i}$ and $w_{j}$ are the joint-embeddings of the two clients. Then, for the given similarity matrix $S$, for each user $i$, we take the top-$N$ in the highest similarity list $\mathcal{S}_{i}$ for each user as the neighbor of user $i$ to build graph $\mathcal{S}^{\prime}$.
4.2.2 Learning global item embedding on the server
Based on this graph, we use a graph-guided aggregation mechanism to update the federated embeddings so that each client can obtain user-specific project embeddings with the help of neighbors with similar preferences. Specifically, the core idea is to use a lightweight Graph Convolution Network (GCN) [7] to update the i-th client item embed by aggregating its neighbors. The convolution operation is as follows:
$\mathcal{R}=\mathcal{S}^{\prime l}A$ (16)
Where A is the round item embedding matrix, the I-th row represents the item embedding obtained from user i, $\mathcal{R}$ is the learned correlation matrix, and the i-th behavior $r_{i}$ In addition, l represents the number of convolution layers, l=1 for simple calculations in this model.
After that, we use the learned correlation matrix $\mathcal{R}$ to sum each user i in an average way to get the global parameter $\theta_{global}$ as (17).
$\theta_{global}=DR$ (17)
Where D represents the degree matrix at the time of aggregation.
4.3 User Regularization
In each round, each user accepts the global parameter $\theta_{global}$ from the server. First of all, the weight of the project embedding layer is initialized with the global shared item embedding $\theta_{global}$. Parameter $\theta_{user,i}$ of the user joint embedding layer, parameter $\theta_{Transformer,i}$ of transformer layer, and fraction prediction function $\theta_{score,i}$ all inherit the model trained in the last round. Then, we will continue to further train the model using the loss function of formula (12) and update the $\theta^{i}_{t}$ with stochastic gradient descent algorithm, and t-th update step can be formulated as follows.
$\theta^{i}_{t}=\theta^{i}_{t}-\eta\nabla_{\theta^{i}_{t}}\mathcal{L}_{all}$ (18)
Where $\eta$ is the learning rate and $\sum_{i=1}^{N}\theta^{i}_{t}\mathcal{L}_{all}$ is the gradient of the model parameters with respect to loss.
4.4 Algorithm
Based on the above, after several rounds of iteration, we can optimize the whole model parameters. The algorithm we propose is shown in Algorithm 1. The whole idea of our algorithm is to use the text feature description of the user to build the user’s relationship graph, and use the relationship graph and federated parameter aggregation method to learn the global item embedding that can reflect the universality.
Algorithm 1 Graph Federation Recommendation System based on User Text description features.
  Initialize $\eta,\lambda,\{\theta_{item,i},\theta_{user,i},\theta_{score,i},\theta_{Transformer,i}\}_{i=1}^{N}$
  for each round $t=1,2,...,T$ do
        Client local training:
            for each client $i=1,2,...,N$ in users do:
                for each local epoch e from 1 to $E$ do:
                    Update($\theta^{(t)}_{item,i},\theta^{(t)}_{user,i},\theta^{(t)}_{score,i},\theta^{(t)}_{Transformer,i}$) with Eq.(18)
                end for
                    ($\theta^{(t+1)}_{item,i},\theta^{(t+1)}_{user,i},\theta^{(t+1)}_{score,i},\theta^{(t+1)}_{Transformer,i}$) $\leftarrow$ ($\theta^{(t)}_{item,i},\theta^{(t)}_{user,i},\theta^{(t)}_{score,i},\theta^{(t)}_{Transformer,i}$)
            end for
        Server update with graph:
                Expand the user joint-embedding layer weight matrix into a vector
                Calculate the matrix of similarity of user text feature descriptions with Eq.(15)
                Build user relationship graph.
                Learn user common item embeddings with Eq.(16)
                Learn globally shared item embedding $\theta_{global}$ with Eq.(17)
4.4.1 Improve client efficiency
In real-world scenarios, user feature attributes and items in recommender systems are often very large, which poses potential embedded storage and communication overhead challenges for resource-limited client devices. To address this problem, we propose that each user is embedded with the user’s textual features only in the first round and when the user’s information changes.
4.5 Privacy Protection
Under the federated learning framework, our approach inherits the privacy benefits of only storing private data locally for each user. However, when the client uploads parameters after each round, there is a certain risk of privacy disclosure. To address this risk, we try to introduce local differential privacy[40] to our method by adding a certain amount of Laplacian noise (be shown at Formula.(17)) to the parameter before uploading it. In doing so, you can improve the effect of privacy protection without losing performance.
$\theta_{i}=\theta_{i}+Laplace(0,\alpha)$ (19)
Where, $\alpha$ is the noise intensity, the larger the $\alpha$, the more noise it introduces, the better the privacy protection performance.
5 Experiment
This section analyzes the proposed methods through experiments, aiming to answer the following questions:
  • 1.
    Q1: Is UFGraphFR superior to current advanced federated and centralized recommendation models?
  • 2.
    Q2: Do the modules we introduce make the unintroduced modules perform better?
  • 3.
    Q3: How will UFGraphFR perform with different pre-trained language models?
  • 4.
    Q4: Does the noise introduced by UFGraphFR in increasing local segmentation privacy affect performance?
5.1 Datasets and Evaluation Protocols
Table 1: Dataset statistics
Name Users Items Interactions Sparsity
MovieLens-100K 943 1682 100,000 93.70%
MovieLens-1M 6,040 3,706 1,000,209 95.53%
Lastfm-2K 1,600 12,454 185,650 99.07%
HetRec2011 2,113 10,109 855,598 95.99%
5.1.1 datasets
We validated the proposed UFGraphFR on four recommended benchmark datasets: MovieLens-100K, MovieLens-1M[41], Lastfm-2K[42] , HetRec2011[42] and Douban [43]. In particular, two MovieLens datasets were collected from the MovieLens website, recording user ratings for movies, with no fewer than 20 ratings per user. Lastfm-2K is a music dataset where each user keeps a list of artists listened to and a listen count. We removed users with less than 5 interactions from Lastfm-2K. HetRec2011 is an extension of MovieLens-10M, which connects movies with the corresponding web pages of the Internet Movie Database (IMDb) and the Rotten Tomatoes movie review system. For the Lastfm-2K and HetRec2011 datasets, we used the user id and the total number of user interaction items as user attributes. Detail statistics are shown in Table.1
5.1.2 Evaluation protocols.
To make a fair comparison, we followed the common leave-one method of evaluating Settings and used hit ratio (HR)[44] and normalized cumulative gain of loss (NDCG)[45] metrics to evaluate performance. The results are expressed in units of 0.01.
5.2 Baselines and Implementation Details
5.2.1 Baselines
We compared our approach to two baseline branches, including a centralized and federated recommendation model. All methods make recommendations based solely on user-project interactions.
  • 1.
    Matrix Factorization (MF)[46] A typical recommendation model. It decomposes the scoring matrix into two embeddings in the same hidden space, describing user and item characteristics respectively.
  • 2.
    Neural Collaborative Filtering (NCF)[47] This method is one of the most representative neural recommendation models. We first learn the user embedded module and the item embedded module, and then model the user’s item interaction by MLP.
  • 3.
    Self-supervised Graph Learning (SGL)[48] This method is a self-supervised graph learning enhanced recommendation model.
  • 4.
    FedMF[4] FedMF trains users locally for embedding and uploads project gradients to the server for global aggregation.
  • 5.
    FedNCF[15] FedNCF treats user embeddings as private components of local training and shares project embeddings and MLPS to perform collaborative training.
  • 6.
    Federated Reconstruction (FedRecon)[14] An advanced personalized federated learning framework, FedRecon retrains user embeddings in each round and computes item gradients based on the retrained user embeddings.
  • 7.
    Meta Matrix Factorization (MetaMF)[13] It is a distributed matrix decomposition framework in which meta-networks are used to generate fractional functional modules and private term embeddings.
  • 8.
    Personalized Federated Recommendation (PFedRec)[16] It is a personalized federated recommendation framework where the server first learns a common project embed for all clients and then fine-tunes the project embed with local data for each client.
  • 9.
    Federated LightGCN (FedLightGCN)[1] The way LightGCN[7] extends to the federated learning framework in paper GPFedRec. In particular, each client trains local LightGCN using a first-order interaction subgraph.
  • 10.
    Federated Graph Neural Network (FedPerGNN)[18] It deploies a graph neural network on each client, and users can integrate high-level user-project information through a graph extension protocol.
  • 11.
    Graph-Guided Personalization for Federated Recommendation (GPFedRec)[1] GPFedRec is a graph-guided federated recommendation system, which proposes a graph-guided aggregation mechanism. Our model is mainly referenced from this model.
5.2.2 Implementation details
We implemented this method based on the Pytorch framework. In addition, we developed a variant of the method based on the original model, named UFGraphFR-Lite. UFGraphFR-Lite can improve operational efficiency by periodically updating user diagrams.
In order to make a fair comparison, we compare performance with GPFedRec and indirectly compare other baseline performances by comparing GPFedRec. Therefore, we set all the hyperparameters of my method to be the same as described in the GPFedRec article. That is, set the embedding size to 32, use a fixed batch size of 256, and set the total training cycle (for centralized methods) or communication cycle (for federated methods) to 100, which allows all methods to converge. Set the number of local training rounds for federated learning to 1. For the fractional function module in our method, we use three hidden layers, MLPs, whose schema is 32→16→8→1. We used MediaPipe[49] to load the USE[50] pre-trained model for text embedding. The text embedding dimension was 100, and the output dimension of the joint embedding layer was 32. At the same time, in order to better capture user details, the user’s mlp layer uses a two-layer mlp layer with a 32-64-32 architecture consistent with the feedforward network architecture of the transformer block we use.
We partitioned the dataset according to the following strategy: the first interaction record of each user as the test set, the second interaction record of each user as the validation set, and the rest of the interaction records of each user (i.e., all records except the first and the second) as the training set. This approach ensures that there is at least one test sample, one validation sample, and multiple training samples for each user, which helps in the evaluation of the model’s generalization ability on different users.
Table 2: Performance comparison on four datasets.
Method MovieLens-100K MovieLens-1M Lastfm-2K HetRec2011
HR@10 NDCG@10 HR@10 NDCG@10 HR@10 NDCG@10 HR@10 NDCG@10
CenRec MF 64.48 38.61 68.69 41.45 83.13 71.78 66.07 41.21
NCF 64.21 37.13 64.02 38.16 82.57 68.26 64.74 39.55
SGL 64.9 40.02 62.6 34.13 82.37 68.59 65.12 40.18
FedRec FedMF 66.17 38.73 67.91 40.81 81.63 68.18 64.69 40.29
FedNCF 60.66 33.93 60.38 34.13 81.44 61.95 60.86 36.27
FedRecon 65.22 38.49 62.78 36.82 82.06 67.37 61.57 34.2
MetaMF 66.21 41.02 44.98 26.31 81.04 64.13 54.52 32.36
PFedRec 71.37 42.59 73.03 44.49 82.38 73.19 67.2 42.7
FedLightGCN 24.53 12.78 37.53 15.01 43.75 15.17 22.65 7.96
FedPerGNN 11.52 5.08 9.31 4.09 10.56 4.25
GPFedRec 72.85 43.77 72.17 43.61 83.44 74.11 69.41 43.34
Ours UFGraphFR 76.03* 47.31* 75.55* 46.32* 85.69* 77.32* 70.42 45.17*
UFGraphFR-Lite 75.19 46.46 71.04 42.49 85.19 77.32 70.56* 44.95
Improvement $\uparrow$4.37 $\uparrow$8.08 $\uparrow$4.68 $\uparrow$6.22 $\uparrow$2.69 $\uparrow$ 4.33 $\uparrow$1.46 $\uparrow$4.22
    Performance comparison of the four datasets. The best results are in bold and the best baseline results are underlined. FedPerGNN could not run on HetRec2011 due to unacceptable memory allocation (indicated by ‘-’). “ to indicate this). “*” and ‘Improvement’ denote statistically significant improvement and performance improvement over the best baseline, respectively. $\uparrow$ denotes the difference in enhancement relative to the best baseline
5.3 Performance (Q1)
Table 2 shows the performance of HR and NDCG on the four data sets in the Top-10 recommended scenarios.
1) Our approach achieves better performance than the centralized recommendation system approach in all Settings. In the Top-10 scenario, the largest performance gains for HR and NDCG were 17.14% and 18.21% on the MovieLens-100K dataset. In a centralized setup, all users share the same item embeddings and scoring functions, with only user embeddings retained for personalization capture. In contrast, our approach uses user embeddings and scoring functions as private components to learn user characteristics. In addition, we introduced graph structures to leverage user embeddings from users to build graphs and mine correlations between clients, which enhanced user preference learning and provided better recommendation performance.
2) Our approach outperforms the federal recommended baseline and achieves state-of-the-art results on almost all datasets. Compared with GPFedRec, our method attempts to use the description of the user’s text features (that is, the structured attributes of the user are natural linguistic and sent into the vector embedded by the pre-trained language model) to build the user relationship graph, which can better capture the relationship between users and achieve better performance. In addition, we tried to introduce a transformer block into the model, which can better capture the sequential relationships between the interaction sequences, which also brings a high-performance improvement.
Finally, our lightweight variant, UFGraphFR-Lite, can reach or even outperform UFGraphFR, striking a good balance between model efficiency and power consumption.UFGraphFR-Lite makes building user-relationship graphs infrequent and can be rebuilt after a few rounds (e.g., 5 rounds, 10 rounds). The UFGraphFR-Lite in Table 2 was experimented with by updating the user-relationship graph every 5 rounds.
5.4 Ablation experiment (Q2)
Table 3 presents the results of ablation experiments on the UFGraphFR model, aiming to evaluate the contribution of two key modules of the model - the Transformer module and the Joint Embedding module - to the recommendation performance under different datasets (MovieLens-100K, MovieLens-1M, Lastfm-2K, HetRec2011). ) module - the Transformer module and the Joint Embedding (JE) module - contribute to the recommendation performance under different datasets (MovieLens-100K, MovieLens-1M, Lastfm-2K, HetRec2011). The table reports the performance of each configuration on both HR@10 and NDCG@10 metrics while giving the improvement values compared to the baseline model GPFedRec.
Table 3: Ablation study results of UFGraphFR on four datasets. GPFedRec is used as the baseline. Performance drops relative to the full model are shown below each variant.
Method MovieLens-100K MovieLens-1M Lastfm-2K HetRec2011
HR@10 NDCG@10 HR@10 NDCG@10 HR@10 NDCG@10 HR@10 NDCG@10
GPFedRec (baseline) 72.85 43.77 72.17 43.61 83.44 74.11 69.41 43.34
UFGraphFR (w/o Transformer) 72.85 44.73 72.80 43.84 83.94 75.66 69.33 43.37
↓ vs Full $\downarrow$3.18 $\downarrow$2.58 $\downarrow$2.75 $\downarrow$2.48 $\downarrow$1.75 $\downarrow$1.66 $\downarrow$1.09 $\downarrow$1.80
$\uparrow$ vs Baseline $\uparrow$0.00 $\uparrow$0.96 $\uparrow$0.63 $\uparrow$0.23 $\uparrow$0.50 $\uparrow$1.55 $\downarrow$0.08 $\uparrow$0.03
UFGraphFR (w/o Joint Embedding) 74.23 46.40 74.40 45.21 85.94 77.52 69.43 42.78
↓ vs Full $\downarrow$1.80 $\downarrow$0.91 $\downarrow$1.15 $\downarrow$1.11 $\uparrow$0.25 $\uparrow$0.20 $\downarrow$0.99 $\downarrow$2.39
$\uparrow$ vs Baseline $\uparrow$1.38 $\uparrow$2.63 $\uparrow$2.23 $\uparrow$1.60 $\uparrow$2.50 $\uparrow$3.41 $\uparrow$0.02 $\downarrow$0.56
UFGraphFR (w/o Both) 73.38 44.69 73.51 44.48 84.50 75.73 69.05 42.39
↓ vs Full $\downarrow$2.65 $\downarrow$2.62 $\downarrow$2.04 $\downarrow$1.84 $\downarrow$1.19 $\downarrow$1.59 $\downarrow$1.37 $\downarrow$2.78
$\uparrow$ vs Baseline $\uparrow$0.53 $\uparrow$0.92 $\uparrow$1.34 $\uparrow$0.87 $\uparrow$1.06 $\uparrow$1.62 $\downarrow$0.36 $\downarrow$0.95
UFGraphFR (Full) 76.03 47.31 75.55 46.32 85.69 77.32 70.42 45.17
    $\downarrow$ denotes the difference in performance degradation relative to the full model of UFGraphFR. $\uparrow$ denotes the difference in enhancement relative to the baseline model GPFedRec. Bolding is the highest performance for the current dataset
As shown in Figure 2, in our experiments, the model that removes the Sequence Modeling module (w/o Transformer) removes the Transformer block layer on the right side; the model that removes the Joint Embedding module (w/o Joint Embedding) removes the left side of the Joint Embedding is replaced with a normal Embedding layer. Therefore, the model that removes both modules (w/o Both) is to implement both of the above implementations simultaneously. However, compared with the baseline model GPFedRecLABEL:zhang_gpfedrec_2024, it still has an additional UserMLP layer for further learning user preferences.
5.4.1 w/o Transformer
With the removal of the Transformer (w/o Transformer), there is a general decrease in performance across the datasets, which is particularly noticeable on NDCG, e.g. MovieLens-1M drops to 43.84, which is a significant decrease compared to the 46.32 of the full model. Transformer better models long-term dependencies and behavioral patterns in user interaction sequences, improving the ranking accuracy of recommendations.
5.4.2 w/o Joint Embedding
On the Lastfm-2K dataset, although the w/o Joint Embedding has some small improvement compared with the full model, which may be due to the failure of the module to play a full role or overfitting. However, the MovieLens series dataset and the HetRec2011 dataset, have a stable improvement on both HR and NDCG, which verifies the universal value of text features in user relationship modeling.
5.4.3 w/o Both
Compared to the benchmark model, there is a small performance gain on most of the datasets, but with some degree of degradation on HetRec2011, which may be related to some of the noise introduced during training on the UserMLP layer. The general decrease in performance compared to the full model suggests that the two modules are complementary to each other and are missing, e.g., on MovieLens-100K, the HR of the full model is 76.03, whereas the w/o Both is only 73.38, a decrease of close to 3 points.
In general, the design of the UFGraphFR model significantly improves the effect of the recommendation system, especially the best performance of the Joint-Embedding and Transformer modules. Ablation experiments show that the complete UFGraphFR model achieves optimal performance through the combination of the two.
5.5 PLM & Modeling Effect(Q3)
In this subsection, we will evaluate the use of different pre-training models to train UFGraphFR. Specifically, we select the following model to train the model on the 100k dataset:
  • 1.
    Universal Sentence Encoder,USE [50] it provided by MediaPipe[49] to embed text. The text embedding dimension is 100.
  • 2.
    LaBSE[51] The language-agnostic BERT sentence embedding encodes text into high-dimensional vectors. The Embedding dimension is 768.
  • 3.
    GTR-T5-Base,T5[52] This is a sentence-transformers model: It maps sentences & paragraphs to a 768 dimensional dense vector space.
  • 4.
    all-MiniLM-L6-v2,MiniLM This is a sentence-transformers model: It maps sentences & paragraphs to a 384 dimensional dense vector space and can be used for tasks like clustering or semantic search.
    TinyBERT-L6,TinyBERT[53] This is a sentence-transformers model: It maps sentences & paragraphs to a 768 dimensional dense vector space and can be used for tasks like clustering or semantic search.
The experimental results are shown in Table 4. It shows the performance comparison of different Pre trained Language Models (PLMs) on recommendation tasks, with evaluation metrics of HR@10 (Hit Rate) and NDCG@10 (Normalized Discounted Cumulative Gain).
The experimental results indicate that all PLMs are HR@10 And NDCG@10 The two indicators are superior to the baseline model GPFedRec, indicating that the graph structure constructed through user text features has significant advantages in federated recommendation. Among all models, TinyBERT performs the best (HR@10 : 76.35, NDCG@10 : 47.36), demonstrating its powerful semantic modeling ability and competitive advantage in lightweight models.
Table 4: Performance comparison of PLMs
Model HR@10 NDCG@10
GPFedRec 72.85 43.77
UFGraphFR-USE 75.72 47.08
UFGraphFR-MiniLM-L6 74.87 46.63
UFGraphFR-T5 76.14 46.60
UFGraphFR-TinyBERT 76.35 47.36
UFGraphFR-LaBSE 75.50 46.89
5.6 Privacy Protection (Q4)
In this subsection, we evaluate the performance of our privacy protection enhanced UFGraphFR with the local differential privacy strategy. Particularly, we set the noise intensity $dp$ = [0.05, 0.1, 0.2, 0.3, 0.4], and experimental results are shown in Table 5.
Table 5: Performance on different dps.
0.05 0.1 0.2 0.3 0.4
HR@10 75.29 75.93 77.20 76.03 75.72
NDCG@10 46.34 46.86 47.62 46.80 47.27
Figure 5: Comparison of model convergence. The horizontal axis is the number of training rounds for federated learning, and the vertical axis is the model performance on both metrics.
When the differential privacy noise ratio is at a low level (dp=0.05 and dp=0.1), the HR and NDCG of the model maintain good performance, reaching 0.7529 and 0.4634 (dp=0.05) and 0.7593 and 0.4686 (dp=0.1), respectively. With dp increasing from 0.05 to 0.1, the HR and NDCG of the model are slightly improved, indicating that at low noise levels, appropriate noise introduction has little impact on model performance, and may even help the generalization ability of the model to avoid overfitting. When dp was increased to 0.2, the model’s HR and NDCG reached 0.7720 and 0.4762, respectively, which were the best results in the test. The results show that the introduction of differential privacy not only effectively protects user privacy, but also optimizes the performance of the model at a moderate noise level. Noise at this level helps the model improve privacy without sacrificing recommendation quality. As dp increases to 0.3 and 0.4, the performance of the model begins to decline. When dp=0.3, HR and NDCG are 0.7603 and 0.4680 respectively, while when dp=0.4, these two indicators fall further to 0.7572 and 0.4727. The further increase of noise leads to the degradation of model performance, which may be because too much noise interferes with the model’s learning of data patterns and weakens the prediction ability of the recommendation system.
The experimental results show that there is an optimal interval (about dp=0.2) in the dp setting of differential privacy noise ratio, which can balance the privacy protection and model performance, protect user privacy, and provide high-quality recommendation results. When dp exceeds this range (e.g. 0.3 or 0.4), noise interference increases, resulting in a decrease in recommendation effectiveness.
5.7 Convergence analysis
We compared the convergence of our method with the baseline GPFedRec that we mainly compared, and there are two main conclusions: First, as shown in Figure 5, both UFGraphFR and UFGraphFR-Lite outperformed BaseLine GPFedRec[1] on HR and NDCG on all data sets, suggesting that the two variants were recommended better than the baseline model. Second, UFGraphFR and UFGraphFR-Lite converge relatively quickly on all datasets, especially within the first 50 Epochs.
In addition, we can see that our method converges quickly on all four datasets and the final result is better than the baseline model GPFedRec. There is less interactive data available for each user modeling preference for the Lastfm-2k dataset. Our approach learns personalized item embedding by aggregating highly similar users using their text feature descriptions, alleviating the difficulty of local personalization modeling and accelerating convergence. Secondly, we introduced a Transformer into the model, which can better model the user interaction sequence to speed up convergence.
6 Conclusion
In this paper, we propose a novel personalized federated recommendation framework, UFGraphFR, which constructs a user relationship graph based on the user’s textual feature descriptions. To protect user privacy, we design a joint embedding layer that transforms locally embedded user features into low-dimensional representations, and utilize the weights of this layer to construct user graphs on the server side without directly exposing raw data.
Furthermore, we integrate a Transformer-based module to capture the temporal dependencies within user-item interaction sequences. Experimental results on multiple benchmark datasets demonstrate that UFGraphFR outperforms existing centralized and federated recommendation models, including the strong baseline GPFedRec, in terms of both HR@10 and NDCG@10 metrics. This validates the effectiveness of utilizing semantic relationships derived from user textual features under the federated learning paradigm.
Despite its superior accuracy, UFGraphFR currently suffers from relatively high computational overhead compared to simpler models such as GPFedRec. This limitation points to future directions where model efficiency can be improved, for example, by incorporating lightweight Transformer variants or dynamic graph updating strategies.
In addition, the current model focuses solely on user-side textual features for semantic enhancement. In future work, we aim to further enrich the model by introducing item-side textual features, enabling bidirectional semantic modeling to further improve recommendation quality under privacy-preserving settings.
7 Declaration of Generative AI
During the preparation of this work the author(s) used Deepseek and ChatGPT in order to polished language. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.

### v4

@@
 Abstract
-Federated learning has become a key paradigm that utilizes the "data available but invisible" feature to protect user privacy, especially in the artificial intelligence field where user privacy is highly valued.
-Federated learning enables users to collaboratively train models without sharing raw data.
-Motivated by this, federated recommendation systems offer a promising architecture that balances user privacy with recommendation accuracy through distributed collaborative learning.
-However, existing federated recommendation methods often neglect the underlying semantic or behavioral relationships between users during parameter aggregation, which limits their recommendation effectiveness.
-To overcome this limitation, graph-based federated recommendation systems have been proposed to leverage neighborhood information.
-Yet, conventional graph construction methods usually require access to raw user data or explicit social links, which contradicts the strict privacy requirements of federated learning.
-In this paper, we propose UFGraphFR, a novel model designed to effectively capture user relationships for constructing user graphs while preserving the confidentiality of users’ raw data.
-(1) To better utilize users’ textual feature descriptions under the hypothesis that users with similar textual profiles exhibit analogous preferences, we propose a Joint-Embedding Layer architecture designed to encode structured user features into fixed-dimensional vector representations.
-(2) To prevent raw data exposure during temporal pattern extraction, we introduce a Transformer-based sequence encoder with differential privacy guarantees, which employs multi-head self-attention mechanisms to model long-range dependencies in user behavioral sequences through on-device processing.
-These embeddings are then used to construct a semantic user graph for server-side aggregation via a lightweight graph neural architecture.
-UFGraphFR ensures privacy-preserving, personalized recommendations while maintaining model robustness and convergence efficiency under heterogeneous conditions.
+Federated learning offers a privacy-preserving framework for recommendation systems by enabling local data processing; however, data localization introduces substantial obstacles.
+Traditional federated recommendation approaches treat each user as an isolated entity, failing to construct global user relationship graphs that capture collaborative signals, which limits the accuracy of recommendations.
+To address this limitation, we derive insight from the insight that semantic similarity reflects preference.
+similarity, which can be used to improve the construction of user relationship graphs.
+This paper proposes UFGraphFR, a novel framework with three key components: 1) On the client side, private structured data is first transformed into text descriptions.
+These descriptions are then encoded into semantic vectors using pre-trained models; 2) On the server side, user relationship graphs are securely reconstructed using aggregated model weights without accessing raw data, followed by information propagation through lightweight graph neural networks; 3) On the client side, user behavior sequences are personalized using Transformer architectures.
+Extensive experiments conducted on four benchmark datasets demonstrate that UFGraphFR significantly outperforms state-of-the-art baselines in both recommendation accuracy and personalization.
+The framework also maintains robustness across different pre-trained models, as evidenced by the consistent performance metrics obtained.
+This work provides a practical method for efficient federated recommendations with strict privacy by using semantic vectors, secure user relationship graphs, and personalized behavior sequences.
 The code is available at: https://github.com/trueWangSyutung/UFGraphFR.
 keywords:
-Federated Learning;Federated Recommendation System;Recommendation
-journal: Arxiv
-[…]
-[label1]organization=School of Computer Science and Engineering, addressline=Tianjin University of Technology, city=Tianjin, postcode=300384, state=Tianjin, country=China
-{highlights}
-Proposed a joint-embedding layer that converts structured user data into privacy-preserving text-based representations.
-Developed a differentially private semantic graph construction method to model user correlations in federated learning.
-Demonstrated superior performance over both federated and centralized baselines across multiple real-world datasets.
-Validated framework effectiveness with different pre-trained language model embeddings.
+Federated Learning, Federated Recommendation System, Recommendation
 1 Introduction
-In the era of big data and privacy computing, intelligent recommendation systems have become indispensable infrastructure for modern digital services, enabling personalized movie suggestions, targeted food delivery recommendations, and even emotion-aware content curation through user comment analysis[1].
-However, the conventional centralized paradigm that aggregates raw user data on servers poses critical privacy risks, as evidenced by growing user reluctance to share personal information with service providers[2].
-In recent years, with the proposal of federated learning, which is a privacy-preserving framework that aims to better protect users’ privacy[3].
-More and more research directions have started to apply this framework.
-Its emergence has catalyzed its application in the field of recommender systems, allowing the training of complete recommender system models without the raw data leaving the client device locally[4] .
-Despite significant progress, there are two fundamental limitations of existing joint recommendation methods: (1).
-They model users as isolated entities mainly through local training goals, ignoring the inherent social correlations and behavioral dependencies among user groups[5], such as belonging to the same Beijing area, belonging to the same male gender, etc.
-This oversimplification contradicts real-world observations, where users with common interests often exhibit collaborative consumption patterns.
-(2).
-Despite the remarkable success of graph-enhanced recommender systems in centralized environments through neighborhood information aggregation, their direct application in federated learning is still prohibited due to privacy violations resulting from explicit user relationship disclosure[6] Combining the problems of the above two points, a graph augmentation based federated recommender system is proposed [7, 8].
-(3).
-Current approaches mainly rely on Users’ ID-based feature representations[9], and such recommender systems perform poorly in the cold-start problem of models.
-Therefore, semantics can be utilized to enhance the cold start of the model.
-For example, two users can establish an association by the similarity between their textual features.
-In this paper, we propose a novel framework for personalized federated recommendation, termed UFGraphFR (User Text description-based Feature Graph Federated Recommendation), which constructs a user relationship graph based on structured textual features without exposing private data.
-The proposed framework enables the server to aggregate user embeddings and generate personalized recommendations while complying with the privacy-preserving principle of federated learning.
-Specifically, our main contributions are summarized as follows:
-  • 1.
-    To better utilize user-structured attribute data, we introduce a joint-embedding layer that transforms user-side structured features into natural language descriptions, encodes them via a pre-trained language model into high-dimensional embeddings, and further projects them into compact representations via a lightweight linear transformation.
-The projection weights are shared with the server to guide semantic graph formation.
-  • 2.
-    To address the limitation that existing federated recommender systems ignore correlations between users, we propose a user graph construction method that builds a semantic user relationship graph on the server using differentially private embedding weights of user text descriptions.
-This allows the system to model collaborative preferences while maintaining strict data privacy.
-  • 3.
-    To validate the effectiveness of the proposed framework, we conduct extensive experiments on multiple real-world datasets.
-Results show that UFGraphFR achieves superior performance compared to existing federated and centralized baselines while preserving user privacy.
-And verify the effectiveness under different pre-trained language model embeddings.
+In the era of big data and privacy computing, intelligent recommendation systems have become an indispensable infrastructure for modern digital services, enabling personalized movie suggestions, targeted food delivery recommendations, and even emotion-aware content curation through user comment analysis[1].
+However, conventional centralized paradigms that aggregate raw user data on servers pose critical privacy risks, as evidenced by growing user reluctance to share personal information with service providers[2].
+Federated learning emerges as a promising privacy-preserving framework, allowing collaborative model training without exposing users’ private data [3].
+Its application in recommender systems enables the construction of complete models while keeping raw data localized on client devices[4].
+However, a fundamental limitation of existing federated recommendation approaches stems from the isolation of raw user-item interaction records on local devices, which prevents the server from accessing the global user-item interaction matrix[5, 6].
+Consequently, unlike traditional graph neural network recommendation models that leverage global graph structures (such as user-item bipartite graphs or user social graphs) to capture complex user correlations and collaborative signals, federated systems primarily rely on limited local data from individual users.
+This constraint inherently caps the upper bound of recommendation accuracy, as models cannot effectively exploit the rich relational patterns available in centralized settings.
+In this paper, we propose a novel framework for personalized federated recommendation, termed UFGraphFR (User text-description Feature Graph Federated Rec), which addresses a key challenge: how to find a ’proxy’ signal that enables the server to measure user similarity and approximately build a ’user relationship graph’ without exposing raw private data.
+To solve this problem, UFGraphFR securely transforms local private structured attributes into textual descriptions and maps them into a unified semantic space using pre-trained language models, thereby enabling safe user relationship construction.
+The main contributions of this work are summarized as follows:
+  • We innovatively leverage user text descriptions as a secure bridge for constructing user relation graphs, breaking away from ID-centric paradigms and enabling semantic-aware similarity modeling under privacy constraints.
+  • We design a learning architecture: on the client side, private structured data is first converted into text descriptions and encoded into semantic vectors using pretrained models; on the server side, user relation graphs are reconstructed securely using aggregated model weights without accessing raw data, followed by information propagation via lightweight graph neural networks.
+  • We propose the complete UFGraphFR framework, which further incorporates Transformer architectures on the client side to capture personalized user behavior sequences.
+  • Through comprehensive experiments on four benchmark datasets, we validate the superiority of UFGraphFR in recommendation accuracy, personalization capability, and robustness, supported by in-depth ablation studies.
+The rest of this paper is organized as follows: Section II reviews the related work on federated recommender systems and graph-based learning recommender systems.
+Section III introduces the UFGraphFR method proposed in this paper, including its overall framework and key technical details.
+Section IV will focus on the experimental part, including performance experiments and ablation experiments.
 2 Related Work
+Table 1: Comparison of related work.
+Method Core Idea Pros Cons/Limitations
+FedMF [4] Applies matrix factorization within the FL framework.
+Users train locally and upload gradients.
+Simple and efficient; Establishes a baseline for FL in recommendation.
+Ignores user-user correlations; Limited by the linear nature of MF.
+FedNCF [7] Extends Neural Collaborative Filtering to FL, using MLPs for non-linear user-item interaction modeling.
+Captures non-linear interactions; More expressive than linear models like FedMF.
+Still treats users in isolation; High communication cost due to large MLP parameters.
+PFedRec [8] Removes user embeddings from the server, learning personalized scoring functions on each client.
+Achieves high personalization; Decouples user and item learning.
+Fails to model collaborative signals between similar users; Performance may degrade for users with sparse data.
+FedRecon [9] Retrains user embeddings in every communication round based on local data.
+Enhances personalization by avoiding stale embeddings.
+Computationally expensive for clients; May lead to unstable convergence.
+FedPerGNN [10] Deploys GNNs on clients and uses a graph extension protocol to find higher-order neighbors.
+Leverages structural information for better recommendations.
+Requires aligning historical interactions with the server, risking privacy leaks; High computational and communication overhead.
+GPFedRec [1] Proposes a graph-guided aggregation mechanism to capture user preference correlations.
+Explicitly models user-user relationships; Improves performance by leveraging neighborhood information.
+The method for constructing the user graph is not explicitly detailed in the provided text, potentially relying on sensitive data.
+EasyRec [11] A centralized system using a text-behavior alignment framework with contrastive learning and language model tuning.
+Effectively integrates rich semantic information from text; Strong generalization capability.
+Centralized architecture violates user privacy; Not applicable to privacy-sensitive federated settings.
+  • This table highlights the trade-offs between personalization, privacy, and the ability to model user correlations in current approaches, motivating the need for a method like UFGraphFR that addresses these limitations
 2.1 Federated Recommendation Systems
 Federated Recommendation Systems (FedRS, as shown in Fig.1) is an important extension of Federated Learning (FL) in the field of recommendation, which realizes cross-client personalization modeling by protecting the user’s privacy.
-Personalized modeling, allows the user’s raw data to complete the training process of the recommender system without leaving the client while protecting the user’s privacy.
-Some existing basic recommender system approaches (e.g., Matrix-Factorization-Based Recommender Systems[10], Neural-Collaborative-Filtering-Based Recommender Systems[11]) have been combined with the FL framework to form, for example, matrix factorization-based federated recommender systems FCF[12], FedMF[4], MetaMF[13], FedRecon[14], and neural collaborative filtering-based federated recommender system FedNCF[15].After this, more and more FedRSs were proposed.
+Personalized modeling, allows the user’s raw data to complete the training process of the Recommender system(RecSys) without leaving the client while protecting the user’s privacy.
+Some existing basic recommender system approaches (e.g., Matrix-Factorization-Based RecSys[12], Neural Collaborative Filtering Based RecSys[13]) have been combined with the FL framework to form, for example, matrix factorization based federated recommender systems FCF[14], FedMF[4], MetaMF[15], FedRecon[9], and neural collaborative filtering based federated recommender system FedNCF[7].
+After this, more and more FedRSs were proposed.
 Zhang et al.[8] proposed a personalized FedRec framework, PFedRec, which removes user embedding and learns personalized scoring functions to capture user preferences.
@@
 As a result, federal recommendations are gradually emerging, which combine privacy protection technologies to protect users’ private data from being leaked.
+To provide a systematic overview of the existing federated recommendation landscape, Table 1 summarizes and compares the key characteristics, advantages, and limitations of several representative methods discussed in this section.
 3 Preliminary
+Table 2: Summary of key notations.
+Symbol Definition
+$U,N$ Set and number of users.
+II Set of items.
+$r_{ui}$ Interaction between user $u$ and item $i$.
+$f(\theta)$ Recommendation model with parameters $\theta$.
+$\theta_{\text{user}},\theta_{\text{item}}$ User and item embedding parameters.
+$\theta_{\text{global}},\theta_{\text{local}}$ Global shared and local private model parameters.
+$\mathcal{G(U,E)}$ User relationship graph with adjacency matrix $\mathcal{A}$.
+$\mathcal{L}_{i}(\theta)$ Local loss on client $i$.
+$\omega_{i}$ Client $i$’s aggregation weight.
+$P_{u},v_{u}$ User $u$’s attribute and PLM-generated text embedding.
+$e_{u}$ Final low-dimensional user embedding.
+$d,d_{1},d_{2}$ Dimensions: joint embedding, PLM output, linear layer output.
+$L$ Length of user’s interaction sequence.
+$e_{i_{t}}$ Embedding of the $t$-th interacted item.
+$d_{\text{item}}$ Item embedding dimension.
+$K,d_{h}$ Number and size of Transformer attention heads.
+$\eta,\lambda$ Learning rate and regularization coefficient.
+$\alpha$ Noise scale for differential privacy.
 Federated Recommendation.
@@
 where $\mathcal{L}_{i}(\theta)$ is the loss of the local client participating in the training, and the parameter $\theta^{*}$ is learned by minimizing the local loss of all clients with the client weight $\omega_{i}$.
+For clarity, we summarize the key notations used throughout this paper in Table 2. These symbols are fundamental to the mathematical formulation and description of the proposed UFGraphFR framework.
 4 Methodology
 In this section, we present the Graph Federation Recommendation System based on User Text description features(UFGraphFR). There are four steps in each communication round:
-  • 1.
   • Local Training: Clients initialize project embeddings using global parameters and train recommendation models with private interaction data.
-  • 2.
   • Parameter Uploading: Clients transmit user joint embedding weights and local item embeddings to the server.
-  • 3.
   • Graph Aggregation: The server constructs user relation graphs from text embeddings and aggregates parameters through graph convolution.
-  • 4.
   • Global Distribution: Updated global item embeddings are broadcast to all clients for next-round initialization.
+In the following sections, we will provide a detailed introduction to the proposed UFGraphFR, as illustrated in Figure 2.
+Figure 2: This figure shows a recommendation system framework that integrates federated learning and graph neural networks.
+Local training of users and item embeddings on the client side, uploading model parameters to the server; The server constructs a relationship graph based on user similarity, aggregates and generates global item embeddings, and sends them back to the client to achieve collaborative recommendation under privacy protection.
 4.1 Local Training
 As shown in Figure 3, the client recommendation model consists of five major modules: 1) joint embedding layer (module parameter $\theta_{user}$) maps user structured attributes to vector space through natural language prompts templates, using a combination of a pre-trained language model initialization and a trainable linear layer; 2) item embedding layer (module parameter $\theta_{item}$) encodes user interaction item IDs as dense vectors; 3), textbfTemporal Transformer module (module parameter $\theta_{Transformer}$) models long-term interaction sequences through the mechanism of multi-head self-attention dependencies; 4) User Feature Refinement MLP (module parameter is $\theta_{\text{umlp}}$) extracts higher-order user representations through a three-layer fully-connected network; and 5) Predictive Scoring Function (module parameter is $\theta_{\text{score}}$) computes user-item embedding splicing based on final prediction results.
-Figure 2: Client Model:The image shows our client model, with the yellow box on the left indicating the structure of the joint embedding layer (consisting of the prompt layer, pre-trained language model layer, and linear layer used to convert structured attributes into natural language).
-On the right is the Transformer Block used for sequence modeling
 4.1.1 Joint Embedding Layer
@@
 These are the Prompts Output layer, the untrainable pre-trained language model (PLM) layer, and the trainable output layer.
-Prompts Output Layer: Each user $u$, has a structured data $P_{u}$, as shown in Fig. 3. The role of Prompts Output Layer is to transform the structured data into textual language.
-Figure 3: Prompts Output Layer: Structured data of the user on the left, textual description on the right
+Prompts Output: Each user $u$, has a structured data $P_{u}$. The role of Prompts Output is to transform the structured data into textual language.
 PLM Layer: After converting the structured features $P_{u}$ into natural language cues via $\Gamma(\dots)$, we obtain a string of user text features.
@@
 where $e_{u}$ is the vector embedded by the joint embedding layer, $W_{d_{1}\times d}$ is the weight matrix of this linear layer and $b$ is the bias of this linear layer.
+Figure 3: Client Model:The image shows our client model, with the yellow box on the left indicating the structure of the joint embedding layer (consisting of the pre-trained language model layer, and linear layer).
+On the right is the Transformer Block used for sequence modeling.
 4.1.2 Transformer Block
@@
 $e_{i_{t}}=E_{\text{item}}(i_{t})\in\mathbb{R}^{d_{\text{item}}}\quad(1\leq t\leq L)$ (4)
-These embeddings are concatenated into a sequence matrix $X\in\mathbb{R}^{L\times d_{\text{item}}}$, which serves as the input to the Transformer Block.
+[…]
 Multi-Head Self-Attention Mechanism To capture the complex dependencies within user interaction sequences, we adopt a multi-head self-attention mechanism.
-Specifically, for each item embedding in the sequence, we compute contextualized representations by projecting the input matrix $X\in\mathbb{R}^{L\times d_{\text{item}}}$ into query ($Q$), key ($K$), and value ($V$) spaces through trainable linear transformations:
+Specifically, for each element embeddeded in the sequence, we compute contextualized representations by projecting the input matrix $X\in\mathbb{R}^{L\times d_{\text{item}}}$ into theto the query ($Q$),KK)($K$) and value ($V$) spaces through trainable linear transformations:
 $Q=XW^{Q}\quad(W^{Q}\in\mathbb{R}^{d_{\text{item}}\times d_{h}})$ (5)
@@
 The first part is the binary cross entropy Loss (BCE Loss) of prediction score and label, as shown in formula (10).
-$\mathcal{L}_{1}(y,\hat{y})=-(ylog(\hat{y})+(1-y)log(1-\hat{y}))$ (10)
+$)\mathcal{L}_{1}(y,\hat{y})=-(ylog(\hat{y})+(1-y)log(1-\hat{y}$)) (10)
 where $y$ represents the user’s label, and $\hat{y}$ represents the predicted score output by the model.
@@
 4.2 Federated Training
-Figure 4: Service Model: First, it receives the parameters passed up from the client; second, it constructs the user relationship graph based on the parameters of the Jointed-Embedding layer; third, it learns the global item embedding; finally, it sends the global item embedding down to the client.
 In the federated learning scenario, which is shown in Fig.
-4, where we classify the parameters of the user model into two categories, and the model parameters are classified into two categories based on privacy sensitivity: globally shared parameters $\theta_{global}$ and local private parameters $\theta_{local}$.
+2, where we classify the parameters of the user model into two categories, and the model parameters are classified into two categories based on privacy sensitivity: globally shared parameters $\theta_{global}$ and local private parameters $\theta_{local}$.
 At each time, the user only uploads the globally shared parameters to the server for parameter aggregation.
@@
 Step2: Similarity computation constructed graph In the second step, we use cosine similarity as a similarity measure between two vectors $w_{i}$ and $w_{j}$ for users $i$ and $j$, and the similarity between customer $i$ and $j$ can be expressed as (15).
-$\mathcal{S}_{ij}=\frac{w_{i}\textperiodcentered w_{j}}{||\vec{w}_{i}||||\vec{w}_{j}||}$ (15)
+[…]
 where $w_{i}$ and $w_{j}$ are the joint-embeddings of the two clients.
@@
 $\mathcal{R}=\mathcal{S}^{\prime l}A$ (16)
-Where A is the round item embedding matrix, the I-th row represents the item embedding obtained from user i, $\mathcal{R}$ is the learned correlation matrix, and the i-th behavior $r_{i}$ In addition, l represents the number of convolution layers, l=1 for simple calculations in this model.
+where A is the round item embedding matrix, the I-th row represents the item embedding obtained from user i, $\mathcal{R}$ is the learned correlation matrix, and the i-th behavior $r_{i}$ .
+In addition, l represents the number of convolution layers, l=1 for simple calculations in this model.
 After that, we use the learned correlation matrix $\mathcal{R}$ to sum each user i in an average way to get the global parameter $\theta_{global}$ as (17).
@@
 The whole idea of our algorithm is to use the text feature description of the user to build the user’s relationship graph, and use the relationship graph and federated parameter aggregation method to learn the global item embedding that can reflect the universality.
-Algorithm 1 Graph Federation Recommendation System based on User Text description features.
-  Initialize $\eta,\lambda,\{\theta_{item,i},\theta_{user,i},\theta_{score,i},\theta_{Transformer,i}\}_{i=1}^{N}$
-  for each round $t=1,2,...,T$ do
-        Client local training:
-            for each client $i=1,2,...,N$ in users do:
-                for each local epoch e from 1 to $E$ do:
-                    Update($\theta^{(t)}_{item,i},\theta^{(t)}_{user,i},\theta^{(t)}_{score,i},\theta^{(t)}_{Transformer,i}$) with Eq.(18)
-                end for
-                    ($\theta^{(t+1)}_{item,i},\theta^{(t+1)}_{user,i},\theta^{(t+1)}_{score,i},\theta^{(t+1)}_{Transformer,i}$) $\leftarrow$ ($\theta^{(t)}_{item,i},\theta^{(t)}_{user,i},\theta^{(t)}_{score,i},\theta^{(t)}_{Transformer,i}$)
-            end for
-        Server update with graph:
-                Expand the user joint-embedding layer weight matrix into a vector
-                Calculate the matrix of similarity of user text feature descriptions with Eq.(15)
-                Build user relationship graph.
-                Learn user common item embeddings with Eq.(16)
-                Learn globally shared item embedding $\theta_{global}$ with Eq.(17)
 4.4.1 Improve client efficiency
@@
 $\theta_{i}=\theta_{i}+Laplace(0,\alpha)$ (19)
-Where, $\alpha$ is the noise intensity, the larger the $\alpha$, the more noise it introduces, the better the privacy protection performance.
+where $\alpha$ is the noise intensity, the larger the $\alpha$, the more noise it introduces, the better the privacy protection performance.
 5 Experiment
 This section analyzes the proposed methods through experiments, aiming to answer the following questions:
-  • 1.
   • Q1: Is UFGraphFR superior to current advanced federated and centralized recommendation models?
-  • 2.
   • Q2: Do the modules we introduce make the unintroduced modules perform better?
-  • 3.
   • Q3: How will UFGraphFR perform with different pre-trained language models?
-  • 4.
   • Q4: Does the noise introduced by UFGraphFR in increasing local segmentation privacy affect performance?
+Algorithm 1 The Proposed UFGraphFR Algorithm
+0: Number of clients NN, total communication rounds $T$, local epochs $E$, learning rate $\eta$, regularization coefficient $\lambda$, user joint-embedding layer parameters $\theta_{user}$, item embedding layer parameters $\theta_{item}$, and other local model parameters $\theta_{local}$.
+0: Optimized global item embedding $\theta_{global}$.
+1:Initialize global parameters $\theta_{global}^{(0)}$ (e.g., item embeddings) on the server.
+2:Broadcast $\theta_{global}^{(0)}$ to all NN clients.
+3:for each communication round t=1,2,,Tt=1,2,...,T do
+4:  Client Local Training (Parallel Execution):
+5:  for each client i=1,2,,Ni=1,2,...,N do
+6:   Initialize local item embedding $\theta_{item,i}^{(t)}\leftarrow\theta_{global}^{(t-1)}$.
+7:   Use local data to train the local model for $E$ epochs by minimizing the loss function $\mathcal{L}_{all}$ (Eq. (12)) via SGD (Eq. (18)).
+8:   Obtain updated local parameters: $\{\theta_{item,i}^{(t)},\theta_{user,i}^{(t)},\theta_{Transformer,i}^{(t)},\theta_{umlp,i}^{(t)},\theta_{score,i}^{(t)}\}$.
+9:   Upload the user joint-embedding weight matrix $W_{i}^{(t)}\in\mathbb{R}^{d_{1}\times d_{2}}$ and the local item embedding $\theta_{item,i}^{(t)}$ to the server.
+10:  end for
+11:  Server Update with Graph Aggregation:
+12:  Vectorize: Flatten each client’s uploaded joint-embedding weights: $w_{i}^{(t)}=\text{vec}(W_{i}^{(t)})\in\mathbb{R}^{d_{1}d_{2}}$.
+13:  Compute Similarity: Calculate cosine similarities $\mathcal{S}_{ij}^{(t)}$ between all $w_{i}^{(t)}$ and $w_{j}^{(t)}$.
+14:  Build Graph: Construct user graph $\mathcal{G}^{(t)}$ by connecting each user to its top-$k$ similar users. Obtain adjacency matrix $\mathcal{A}^{(t)}$ and degree matrix $\mathcal{D}^{(t)}$.
+15:  Aggregate Embeddings: Apply GCN to local item embeddings $A^{(t)}$: $\mathcal{R}^{(t)}=(\mathcal{D}^{(t)})^{-1/2}\mathcal{A}^{(t)}(\mathcal{D}^{(t)})^{-1/2}A^{(t)}$.
+16:  Update Global Embedding: Set $\theta_{\text{global}}^{(t)}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{R}_{i}^{(t)}$.
+17:  6. Broadcast $\theta_{\text{global}}^{(t)}$ to all clients.
+18:end for
+19:return The final global item embedding $\theta_{global}^{(T)}$.
 5.1 Datasets and Evaluation Protocols
@@
 5.1.1 datasets
-We validated the proposed UFGraphFR on four recommended benchmark datasets: MovieLens-100K, MovieLens-1M[41], Lastfm-2K[42] , HetRec2011[42] and Douban [43].
+We validated the proposed UFGraphFR on four recommended benchmark datasets: MovieLens-100K, MovieLens-1M[37], Lastfm-2K[38] and HetRec2011[38].
 In particular, two MovieLens datasets were collected from the MovieLens website, recording user ratings for movies, with no fewer than 20 ratings per user.
@@
 We compared our approach to two baseline branches, including a centralized and federated recommendation model. All methods make recommendations based solely on user-project interactions.
-  • 1.
-    Matrix Factorization (MF)[46] A typical recommendation model. It decomposes the scoring matrix into two embeddings in the same hidden space, describing user and item characteristics respectively.
-  • 2.
+  • Matrix Factorization (MF)[41] A typical recommendation model. It decomposes the scoring matrix into two embeddings in the same hidden space, describing user and item characteristics, respectively.
   • Neural Collaborative Filtering (NCF)[42] This method is one of the most representative neural recommendation models.
-We first learn the user embedded module and the item embedded module, and then model the user’s item interaction by MLP.
-  • 3.
+We first learn the user-embedded module and the item-embedded module, and then model the user’s item interaction by MLP.
   • Self-supervised Graph Learning (SGL)[43] This method is a self-supervised graph learning enhanced recommendation model.
-  • 4.
-    FedMF[4] FedMF trains users locally for embedding and uploads project gradients to the server for global aggregation.
-  • 5.
+  • FedMF[4] FedMF trains users locally to embed and uploads project gradients to the server for global aggregation.
   • FedNCF[7] FedNCF treats user embeddings as private components of local training and shares project embeddings and MLPS to perform collaborative training.
-  • 6.
   • Federated Reconstruction (FedRecon)[9] An advanced personalized federated learning framework, FedRecon retrains user embeddings in each round and computes item gradients based on the retrained user embeddings.
-  • 7.
   • Meta Matrix Factorization (MetaMF)[15] It is a distributed matrix decomposition framework in which meta-networks are used to generate fractional functional modules and private term embeddings.
-  • 8.
   • Personalized Federated Recommendation (PFedRec)[8] It is a personalized federated recommendation framework where the server first learns a common project embed for all clients and then fine-tunes the project embed with local data for each client.
-  • 9.
-    Federated LightGCN (FedLightGCN)[1] The way LightGCN[7] extends to the federated learning framework in paper GPFedRec.
+  • Federated LightGCN (FedLightGCN)[1] The way LightGCN[5] extends to the federated learning framework in the GPFedRec paper.
 In particular, each client trains local LightGCN using a first-order interaction subgraph.
-  • 10.
-    Federated Graph Neural Network (FedPerGNN)[18] It deploies a graph neural network on each client, and users can integrate high-level user-project information through a graph extension protocol.
-  • 11.
+  • Federated Graph Neural Network (FedPerGNN) [10] It deploys a graph neural network on each client, and users can integrate high-level user-project information through a graph extension protocol.
   • Graph-Guided Personalization for Federated Recommendation (GPFedRec)[1] GPFedRec is a graph-guided federated recommendation system, which proposes a graph-guided aggregation mechanism.
@@
 5.2.2 Implementation details
-We implemented this method based on the Pytorch framework.
-In addition, we developed a variant of the method based on the original model, named UFGraphFR-Lite.
+We implemented this method on the basis of the Pytorch framework.
+In addition, we developed a variant of the original model-based method, named UFGraphFR-Lite.
 UFGraphFR-Lite can improve operational efficiency by periodically updating user diagrams.
@@
 Table 4: Performance comparison on four datasets.
+00footnotetext: Performance comparison of the four datasets.
+The best results are in bold and the best baseline results are underlined.
+FedPerGNN could not run on HetRec2011 due to unacceptable memory allocation (indicated by ‘-’).
+“ to indicate this).
+“*” and ‘Improvement’ denote statistically significant improvement and performance improvement over the best baseline, respectively.
+$\uparrow$ denotes the difference in enhancement relative to the best baseline
 Method MovieLens-100K MovieLens-1M Lastfm-2K HetRec2011
@@
 Improvement $\uparrow$4.37 $\uparrow$8.08 $\uparrow$4.68 $\uparrow$6.22 $\uparrow$2.69 $\uparrow$ 4.33 $\uparrow$1.46 $\uparrow$4.22
-    Performance comparison of the four datasets.
-The best results are in bold and the best baseline results are underlined.
-FedPerGNN could not run on HetRec2011 due to unacceptable memory allocation (indicated by ‘-’).
-“ to indicate this).
-“*” and ‘Improvement’ denote statistically significant improvement and performance improvement over the best baseline, respectively.
-$\uparrow$ denotes the difference in enhancement relative to the best baseline
 5.3 Performance (Q1)
@@
 In addition, we tried to introduce a transformer block into the model, which can better capture the sequential relationships between the interaction sequences, which also brings a high-performance improvement.
-Finally, our lightweight variant, UFGraphFR-Lite, can reach or even outperform UFGraphFR, striking a good balance between model efficiency and power consumption.UFGraphFR-Lite makes building user-relationship graphs infrequent and can be rebuilt after a few rounds (e.g., 5 rounds, 10 rounds).
+Finally, our lightweight variant, UFGraphFR-Lite, can reach or even outperform UFGraphFR, striking a good balance between model efficiency and power consumption.
+UFGraphFR-Lite makes building user-relationship graphs infrequent and can be rebuilt after a few rounds (e.g., 5 rounds, 10 rounds).
 The UFGraphFR-Lite in Table 4 was experimented with by updating the user-relationship graph every 5 rounds.
@@
 The table reports the performance of each configuration on both HR@10 and NDCG@10 metrics while giving the improvement values compared to the baseline model GPFedRec.
-Table 3: Ablation study results of UFGraphFR on four datasets. GPFedRec is used as the baseline. Performance drops relative to the full model are shown below each variant.
+Table 5: Ablation study results of UFGraphFR on four datasets.
+GPFedRec is used as the baseline.
+Performance drops relative to the full model are shown below each variant.
+00footnotetext: $\downarrow$ denotes the difference in performance degradation relative to the full model of UFGraphFR.
+$\uparrow$ denotes the difference in enhancement relative to the baseline model GPFedRec.
+Bolding is the highest performance for the current dataset.Among them, the variant UFGraphFR-nT(no Transformer) represents the version without Transformer; UFGraphFR-nJ(no Jointed-Embedding) represents the version without Jointed Embedding; UFGraphFR-nB(no Both) represents the version with both removed;
 Method MovieLens-100K MovieLens-1M Lastfm-2K HetRec2011
 HR@10 NDCG@10 HR@10 NDCG@10 HR@10 NDCG@10 HR@10 NDCG@10
-GPFedRec (baseline) 72.85 43.77 72.17 43.61 83.44 74.11 69.41 43.34
-UFGraphFR (w/o Transformer) 72.85 44.73 72.80 43.84 83.94 75.66 69.33 43.37
+GPFedRec 72.85 43.77 72.17 43.61 83.44 74.11 69.41 43.34
+UFGraphFR-nT
+72.85 44.73 72.80 43.84 83.94 75.66 69.33 43.37
 ↓ vs Full $\downarrow$3.18 $\downarrow$2.58 $\downarrow$2.75 $\downarrow$2.48 $\downarrow$1.75 $\downarrow$1.66 $\downarrow$1.09 $\downarrow$1.80
 $\uparrow$ vs Baseline $\uparrow$0.00 $\uparrow$0.96 $\uparrow$0.63 $\uparrow$0.23 $\uparrow$0.50 $\uparrow$1.55 $\downarrow$0.08 $\uparrow$0.03
-UFGraphFR (w/o Joint Embedding) 74.23 46.40 74.40 45.21 85.94 77.52 69.43 42.78
+UFGraphFR-nJ
+74.23 46.40 74.40 45.21 85.94 77.52 69.43 42.78
 ↓ vs Full $\downarrow$1.80 $\downarrow$0.91 $\downarrow$1.15 $\downarrow$1.11 $\uparrow$0.25 $\uparrow$0.20 $\downarrow$0.99 $\downarrow$2.39
 $\uparrow$ vs Baseline $\uparrow$1.38 $\uparrow$2.63 $\uparrow$2.23 $\uparrow$1.60 $\uparrow$2.50 $\uparrow$3.41 $\uparrow$0.02 $\downarrow$0.56
-UFGraphFR (w/o Both) 73.38 44.69 73.51 44.48 84.50 75.73 69.05 42.39
+UFGraphFR-nB
+73.38 44.69 73.51 44.48 84.50 75.73 69.05 42.39
 ↓ vs Full $\downarrow$2.65 $\downarrow$2.62 $\downarrow$2.04 $\downarrow$1.84 $\downarrow$1.19 $\downarrow$1.59 $\downarrow$1.37 $\downarrow$2.78
 $\uparrow$ vs Baseline $\uparrow$0.53 $\uparrow$0.92 $\uparrow$1.34 $\uparrow$0.87 $\uparrow$1.06 $\uparrow$1.62 $\downarrow$0.36 $\downarrow$0.95
-UFGraphFR (Full) 76.03 47.31 75.55 46.32 85.69 77.32 70.42 45.17
-    $\downarrow$ denotes the difference in performance degradation relative to the full model of UFGraphFR.
-$\uparrow$ denotes the difference in enhancement relative to the baseline model GPFedRec.
-Bolding is the highest performance for the current dataset
-As shown in Figure 2, in our experiments, the model that removes the Sequence Modeling module (w/o Transformer) removes the Transformer block layer on the right side; the model that removes the Joint Embedding module (w/o Joint Embedding) removes the left side of the Joint Embedding is replaced with a normal Embedding layer.
-Therefore, the model that removes both modules (w/o Both) is to implement both of the above implementations simultaneously.
-However, compared with the baseline model GPFedRecLABEL:zhang_gpfedrec_2024, it still has an additional UserMLP layer for further learning user preferences.
-5.4.1 w/o Transformer
+UFGraphFR 76.03 47.31 75.55 46.32 85.69 77.32 70.42 45.17
+As shown in Figure 3, in our experiments, the model that removes the Sequence Modeling module (UFGraphFR-nT) removes the Transformer block layer on the right side; the model that removes the Joint Embedding module (UFGraphFR -nJ) removes the left side of the Joint Embedding is replaced with a normal Embedding layer.
+Therefore, the model that removes both modules (UFGraphFR-nB) is to implement both of the above implementations simultaneously.
+However, compared with the baseline model GPFedRec[1], it still has an additional UserMLP layer for further learning user preferences.
+5.4.1 no Transformer
 With the removal of the Transformer (w/o Transformer), there is a general decrease in performance across the datasets, which is particularly noticeable on NDCG, e.g.
@@
 Transformer better models long-term dependencies and behavioral patterns in user interaction sequences, improving the ranking accuracy of recommendations.
-5.4.2 w/o Joint Embedding
+5.4.2 no Jointed-Embedding
 On the Lastfm-2K dataset, although the w/o Joint Embedding has some small improvement compared with the full model, which may be due to the failure of the module to play a full role or overfitting.
 However, the MovieLens series dataset and the HetRec2011 dataset, have a stable improvement on both HR and NDCG, which verifies the universal value of text features in user relationship modeling.
-5.4.3 w/o Both
+5.4.3 no Both
 Compared to the benchmark model, there is a small performance gain on most of the datasets, but with some degree of degradation on HetRec2011, which may be related to some of the noise introduced during training on the UserMLP layer.
@@
 In this subsection, we will evaluate the use of different pre-training models to train UFGraphFR. Specifically, we select the following model to train the model on the 100k dataset:
-  • 1.
   • Universal Sentence Encoder,USE [45] it provided by MediaPipe[44] to embed text. The text embedding dimension is 100.
-  • 2.
-    LaBSE[51] The language-agnostic BERT sentence embedding encodes text into high-dimensional vectors. The Embedding dimension is 768.
-  • 3.
-    GTR-T5-Base,T5[52] This is a sentence-transformers model: It maps sentences & paragraphs to a 768 dimensional dense vector space.
-  • 4.
-    all-MiniLM-L6-v2,MiniLM This is a sentence-transformers model: It maps sentences & paragraphs to a 384 dimensional dense vector space and can be used for tasks like clustering or semantic search.
+  • LaBSE[46] The language-agnostic BERT sentence embedding encodes text into high-dimensional vectors. TheăEmbedding dimension is 768.
+  • GTR-T5-Base,T5[47] This is a sentence transformers model: It maps sentences & paragraphs to a 768 dimensional dense vector space.
+  • all-MiniLM-L6-v2,MiniLM This is a sentence transformers model: It maps sentences & paragraphs to a 384 dimensional dense vector space and can be used for tasks like clustering or semantic search.
     TinyBERT-L6,TinyBERT[48] This is a sentence-transformers model: It maps sentences & paragraphs to a 768 dimensional dense vector space and can be used for tasks like clustering or semantic search.
@@
 In future work, we aim to further enrich the model by introducing item-side textual features, enabling bidirectional semantic modeling to further improve recommendation quality under privacy-preserving settings.
-7 Declaration of Generative AI
-During the preparation of this work the author(s) used Deepseek and ChatGPT in order to polished language.
-After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.
+Appendix A Previous submission to TNNLS for peer review
+Recommended Decision:
+  • Associate Editor: #1: Reject The overall evaluation of the current version is not so positive.
+  • Associate Editor: #2: Reject The paper proposes a graph-based method for federated learning.
+The reviewers have found strong concerns about the lack of novelty, incomplete evaluations, incomplete references, and unsatisfactory presentations.
+The authors are suggested to substantially improve the work based on the reviewers’ detailed comments.
+Individual Reviews:
+  • Reviewer #1: This paper proposes UFGraphFR, a novel graph-enhanced federated recommendation framework that leverages user text descriptions to build user relationship graphs while ensuring privacy preservation.
+By introducing a joint embedding layer and Transformer-based sequence modeling, UFGraphFR captures user preferences without relying on explicit user interaction history or unique IDs.
+Additionally, a lightweight variant, UFGraphFR-Lite, reduces computational overhead by periodically updating the user graph, striking a balance between efficiency and performance.
+Experimental results on multiple benchmark datasets demonstrate the superior accuracy of UFGraphFR over state-of-the-art federated and centralized models.
+The proposed methods effectively enhance recommendation quality and privacy protection in federated learning scenarios.
+    Pros:
+    • 1
+      The paper clearly outlines the limitations of existing methods and presents a well-defined solution.
+    • 2
+      The innovation level of this paper can be considered moderate to high. It introduces user text features and effectively integrates existing technologies, such as Transformers and GCNs.
+    • 3
+      The structure of this paper is relatively clear and well-organized.
+    Cons:
+    • 1
+      The figure in the text does not seem so clear. For example, some of the elements in Figure 1. are vague.
+    • 2
+      There are some references that are not properly cited. Line 20 of Page 6, the references for FedMF are not shown correctly.
+    • 3
+      The authors should check the typos throughout this paper. Line 56 of Page 3, the word "in" is missing the "i".
+    • 4
+      Although the authors conducted ablation experiments for key components, maybe additional parameter ablation experiments is needed, such as $\lambda$ in the total loss function.
+    • 5
+      UFGraphFR-Lite periodically updates user diagrams, but the specific implementation details for the update frequency or the exact method of periodic updates do not appear to be provided in the paper.
+    • 6
+      The authors are encouraged to provide a detailed complexity analysis and discuss the scalability of the proposed method to large-scale datasets.
+  • Reviewer #2: After going through the paper, although the references are cited accordingly which is pretty impressive the article requires a re-examination in terms of detailed explanation on novelty, the motivation behind in undertaking this research in general along with future scope improvements.
+This article looks a potential candidate for publication and will have be to be rejected if the above criteria are not met.
+  • Reviewer #3: The paper introduces a UFGraphFR framework of Graph Federation Recommendation System based on User Text description Features, which focuses on using the text-attributed graph to enhance the recommendation while protecting the privacy.
+It tackles two main challenges: (1) How to build the graph without privacy exposure.
+(2) How to train the federated model based on the constructed text graph jointly.
+    Strength:
+    • 1
+      The paper addresses a practical problem of federated recommendation with privacy protections, which is a significant topic in the community of federated learning.
+    • 2
+      Abundant Experimental results demonstrate the effectiveness of the proposed method.
+    • 3
+      The paper is readable and make it easy to follow.
+    Weakness:
+    • 1
+      About novelty: Beyond the framework, there seem to be absence of original techniques and methods, with the approach primarily consisting of an amalgamation of existing techniques.
+    • 2
+      About clarity: the paper is filled with many typo errors. For example, the line 56 of page 3 and the line 20 for FedMF of page 6.
+    • 3
+      About experiments: (1) For datasets: It is unfair to compare only on small datasets and more large-scale datasets should be compared.
+(2) For evaluation: how to split the datasets?
+How many clients are participating during the training?
+Whether the data of each client is fairly divided?
+The author needs to provide a detailed explanation.
+Appendix B Previous submission to EAAI for peer review
+Final Decision: Reject
+  • Reviewer #1:
+    The manuscript presents a timely and well-motivated study on privacy-preserving recommendation systems through a federated learning framework.
+The proposed UFGraphFR model effectively combines semantic-aware graph construction with privacy safeguards, particularly through its use of joint-embedding strategies and a Transformer-based sequence encoder with differential privacy.
+The integration of on-device processing and graph neural networks is technically sound and addresses key limitations in prior work.
+However, a few minor points should be considered for improvement.
+The authors are encouraged to clarify the specific architecture and training details of the Joint-Embedding Layer to enhance reproducibility.
+Additionally, while the Transformer-based encoder is a strong component, its computational complexity on edge devices could be briefly discussed.
+A schematic summarizing the overall workflow would aid reader comprehension.
+Lastly, a short discussion on the scalability of UFGraphFR under varying user graph densities would further strengthen the manuscript.
+Overall, the paper is a valuable contribution and merits publication after minor revision.
+  • Reviewer #2: The research paper titled "UFGraphFR: Graph Federation Recommendation System based on User Text description features" discusses a research area of significant interest to the scientific community.
+However, the paper is currently in a suboptimal state and requires a complete overhaul before a final decision can be made.
+Below are few suggestions to improve the quality of the manuscript:
+    • 1
+      It is suggested to briefly discuss the structure of subsequent sections in a single paragraph at the end of introduction section.
+    • 2
+      It is suggested to include a table of notations used in the subsequent analysis.
+    • 3
+      It is suggested to include a table in Section 2 discussing the pros and cons of popular techniques.
+    • 4
+      It is suggested to have the manuscript revised by an expert whose first language is English. Additionally, revise "Where" to lowercase "where" for each equation.
+    • 5
+      It is suggested to include a step-by-step algorithm outlining the proposed model.
+    • 6
+      It is suggested to perform statistical test such as Friedman test, Nemenyi test and Wilcoxon test.
+    • 7
+      It is suggested to include the DOI information for each research paper in the reference section. In addition, include at least three papers from 2025.
+  • Reviewer #4: Rejected Not an interesting paper. just the abstract, literature, and methodology sections.